An object is thrown vertically upwards. It has the speed of 10m/s when it has reached one- half of its maximum height
a/ how high it rise
b/what is its velocity and acceleration 1sec after thrown ?
c/ 3 sec after?
d/ what is the average velocity during the first half second?

Grade:11

1 Answers

Antara
2 years ago
Let the object be thrown up with velocity u.
So, max height=u2/2g.
velocity at half max ht=v(say)
So,...v2=u2+2as
     …..v2=u2-2g(u2/4g)
     …..v2=u2-u2/2=u2/2
 so, v=u/rt2 
…...i.e. u=10rt2
  1. max ht=u/2g=100*2/10=20m
  2. After t=1s, velocity=u-gt=10rt2-10*1=10(rt2-1)m/s  and acceleration=g
  3. After t=3s velocity=10rt2-10*3=10(3-rt 2) downwards  and acc=g downwards
  4. displacement after 1st half sec=ut-1/2gt2  =10rt2*1/2-1/2*10*(1/2)2     =5rt2-5/4  so avg vel=disp/time=(5rt2-5/4)/(1/2)=10rt2-5/2  (ans)